Jared earns money by making gifts baskets for Father's Day. He does one filled with fishing supplies and one with a sports fan theme. He makes a profit of $9.00 on the fishing supply basket and a profit of $12.50 on the sports fan baskets. He only has time to make 100 baskets and needs to earn at least $1,000 to cover his summer activites.
Which solution satisfies the system of inequalities?
a)70 fishing baskets and 30 sports fan baskets
b)73 fishing baskets and 27 sports fan baskets
c)72 fishing baskets and 28 sports fan baskets
d)75 fishing baskets and 25 sports fan baskets

Answer by richwmiller(17219)

you can do the math as well as I can.
f*9+12.50*s>=1000
now go through all for answers and see which fits
70*9+30*12.50>=1000,
73*9+12.50*27>=1000,
72*9+12.50*28>=1000,
75*9+12.50*25>=1000
You are lucky today the correct answer is not the last one.